OpenStudy (nerdgamer_):

Add one and one-twelfth plus two and three-twelfths. two and one-third three and one-sixth two and one-sixth three and one-third

OpenStudy (destinymasha):

3 4/12

OpenStudy (destinymasha):

or 3 1/3

OpenStudy (♪chibiterasu):

1 1/12 + 2 3/12 = 3 4/12 4/12 simplifies to 1/3 3 1/3 Answer is D

OpenStudy (nerdgamer_):

DO YOU KNOW WHY ITS THAT THIS SITE says to interact and you need to show me how to get the answer ok

OpenStudy (♪chibiterasu):

Can you like, not just give a direct answer please? >_>

OpenStudy (destinymasha):

they both have .... 12 as their denominator , so all you need to do is add them normally

OpenStudy (♪chibiterasu):

*facepalm*

OpenStudy (destinymasha):

[if you could've just held on i would have;]

OpenStudy (nerdgamer_):

Ok but next time show me or I will have to kick you for not obayingto the rules

OpenStudy (♪chibiterasu):

Only mods can do that lol.

OpenStudy (nerdgamer_):

Show me how to get the answer ok

OpenStudy (destinymasha):

ok , sooooorrrryyyy :}

OpenStudy (nerdgamer_):

I am mod

OpenStudy (nerdgamer_):

Its ok I will let it pass

OpenStudy (♪chibiterasu):

Lol right, then why are you not purple?

OpenStudy (♪chibiterasu):

Reporting isn't something immediate you know. >_>

OpenStudy (destinymasha):

write down the equation, 2 + 1 = 3 , 3/12 + 1/12 = 4/12 , the denominators do not change

OpenStudy (♪chibiterasu):

Yes they do >_>

OpenStudy (nerdgamer_):

So the answer is D Right? thats what I got

OpenStudy (♪chibiterasu):

4/4 = 1 12/4 = 3 1/3 Good job at failing.

OpenStudy (destinymasha):

so.. the answer is 3 and 4/12 ---- 4 divided by 4 = 1, 12 divided by 4 is 3, so 1 and 1/3

OpenStudy (destinymasha):

yes

OpenStudy (nerdgamer_):

Ok thxs

OpenStudy (♪chibiterasu):

Wow I'm surrounded by noobs.